# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ A.2   Proof of Proposition 1
Proposition 1 (a) implies that Type G models yield $\frac{N}{2}$ agglomerations, with further bifurcations of the form $\frac{N}{2}\to\frac{N}{4}\to\frac{N}{8}\to\cdots\to 2\to 1$ expected if $N$ is a power of two (Akamatsu-Takayama-Ikeda-JEDC2012; ikeda2012spatial; Osawa-et-al-JRS2017). For Type L, detailed analyses (e.g., Takayama-Ikeda-Thisse-RSUE2020) confirm the emergence of single-peaked or monocentric patterns (Murata-Thisse-JUE2005; Helpman-Book1998), consistent with Proposition 1 and Fig. 9. Also, AMT-2016 formally compares Forslid-Ottaviano-JoEG2003 and Helpman-Book1998. All available formal results in the literature corroborates with Proposition 1 and the numerical examples presented in Fig. 9. ∎
A.3   Proof of Lemma 4
@@
Under Assumption C, we can choose the same set of eigenvectors for $\mathbf{V}$ and $\mathbf{F}$ because they are both symmetric circulant matrices. Let $\{\bm{z}_{k}\}$ be the set of eigenvectors and let $\omega_{k}$ and $\eta_{k}$ be the eigenvalues of $\mathbf{V}$ and $\mathbf{F}$ associated with $\bm{z}_{k}$, respectively. Then, for each eigenvector $\bm{z}_{k}$ except for $\bm{z}_{0}=\bm{1}$, Eq. A.4 yields
@@
As $\mathbf{F}$ and $\mathbf{V}$ are both symmetric, $\eta_{k}$ and $\omega_{k}$ are both real. Thus, Eq. A.5 implies $\operatorname{sgn}[\eta_{k}]=\operatorname{sgn}[\omega_{k}]$. Therefore, $\bar{\bm{x}}$ is stable under all admissible dynamics if and only if $\omega_{k}<0$ for all $k$, excluding $k=0$ that corresponds to $\bm{z}_{0}=\bm{1}$. Likewise, $\bar{\bm{x}}$ is unstable if and only if $\omega_{k}>0$ for some $k$, again excluding $k=0$.
Suppose exactly one $\omega_{k}$ changes sign from negative to positive at $\phi_{k}^{*}$. Then, from Eq. A.5, the corresponding eigenvalue of the Jacobian matrix of any admissible dynamic at $\bar{\bm{x}}$ must also cross zero at $\phi_{k}^{*}$. Bifurcation theory shows that the system departs from $\bar{\bm{x}}$ along the direction of the associated eigenvector $\bm{z}_{k}$, as it is tangent to the “unstable manifold” at the bifurcation point (see, e.g., Hirsch-et-al-Book2012; Kuznetsov-Book2013).
@@ B.1   Development of high-speed transport networks in Japan
Figure B.1 reports the development of high-speed railway and highway networks, respectively, in Japan between 1970 and 2020. The development of highways and high-speed railway networks in Japan was initiated by the Tokyo Olympics held in 1964. Between 1970 and 2020, the total highway (high-speed railway) length increased from 1,119 km (515 km) to 9,050 km (3,106 km), which is a more than eightfold (sixfold) increase. As both of these networks gradually expanded in the past half century, Japan can be seen as an ideal case to study the comparative statics regarding transport costs.
@@ B.2   Japanese cities and their growths
We identify cities in Japan from the Grid Square Statistics of the Population Census in 1970–2020. A city is defined as an urban agglomeration (UA), which is a set of contiguous 1 km-by-1 km grid cells whose population density is at least 1000/km2 and a total population of at least 10,000. The results are robust for alternative threshold values. Figure B.2 shows the identified 431 UAs in 2020. They occupy 6% of land area while containing 80% of the national population. Populated cells with less than 1,000 residents are shown in grey, with darker grey indicating a larger population. We use only the cells reachable by roads from four major islands (Hokkaido, Honshu, Shikoku, and Kyushu). UAs are identified for each census year during the 1970–2020 period (every five years), and Unique IDs throughout the period are assigned. For the details of the algorithm to construct UAs, see Mori-Murakami-DP2025.
@@
From 1970 to 2020, the total population in these cities grew by $55\%$, while the national population increased by only 21%. The population of the largest city, Tokyo, has grown by 67%, an increase about the same size of the second largest city, Osaka. Figures B.3 and B.6 provide closer looks. Population growth is typically associated with concurrent areal growth (Fig. B.3A), but population density generally decreased during the 50 years period (Fig. B.3B), in particular for small to medium-sized cities, indicating local spreading of these cities. It is noted that the largest cities also experienced local spreading as evident in their spatial distributions shown in Fig. 1C as well as Fig. B.4.
@@ B The population distribution within Nagoya in 1970 and 2020
Figure B.4: Local dispersion of Osaka (rank $=$ 2) and Nagoya (rank $=$ 3).
@@
[…]
@@
Employment distribution. The tendency of local dispersion is observed in alternative indicator of agglomeration other than population. Figure B.6 shows the change in the mean values of the maximum and average employment density within a city across all cities in Japan from 1975 to 2014. Their long-run trend indicates that the geographical distribution of employment in a city has flattened over the past half century.
@@ B.3   Cities in other countries
Figure 2 in Section 1 uses Japanese Census data. To allow a parallel comparison between different countries, we employ the the LandScan™  Global Population Database, developed by the Department of Energy’s Oak Ridge National Laboratory (ORNL), as the basic grid population data to see the evolution of cities. Cities are defined in the same manner as for Japanese cities. To asses compatibility between the LandScan data and the census we compare Census and the LandScan data for Japan. LABEL:fig:Japan(0) and LABEL:fig:Japan(2) are, respectively, based on the Census and the the LandScan data. The LandScan data is only available after 2000, and are only roughly in agreement with the precise data based on Census. It is also noted that, while the LandScan data is available every year, we employ 5-year steps to avoid noises due to its data generation procedure that incorporates various interpolations. Nonetheless, LABEL:fig:gc-ld-other-countries confirms the broad tendency of nationwide concentration and local flattening of the cities. To be consistent with our theory that assume a fixed total population, we normalize the total population in each country to unity. In the cases of France, Germany and Japan, the local flattening is apparent even without the normalization.
@@ Appendix C Eight regions
This section considers agglomeration processes in the $N=8$ circular economy. For a selected model from each model category, we follow stationary equilibria branching from $\bar{\bm{x}}$ and then numerically check the local stability of those stationary equilibrium solutions under the replicator dynamic (Taylor-Jorker-MB1978). See Appendix F for the details of the specific formulations and the parameter settings.
@@
Type L model. Figure C.3 considers a Type L model by Allen-Arkolakis-QJE2014 (Section F.2.4). The model incorporates a local dispersion force but no global dispersion force. The uniform equilibrium $\bar{\bm{x}}$ is stable when transport costs are low (when $\phi$ is close to $1$). If we start from $\bar{\bm{x}}$ and consider the process of a monotonic decrease in $\phi$ from $\phi\approx 1$, then a unimodal pattern emerges due to the bifurcation at ϕsuperscriptitalic-$\phi^{**}$ [Proposition 1 (a)]. This is the bifurcation in the model. When $\phi$ decreases further, the spatial pattern smoothly converges to a full concentration in a single region in the lower extreme ($\phi\approx 0$). The local dispersion force is less important than the benefits of agglomeration when interregional transportation is prohibitively costly. Mobile agents prefer concentrating on a smaller number of regions because of the agglomeration forces. As $\phi$ increases, agglomeration force due to costly transportation diminishes, and the relative rise in the local dispersion force induces a crowding-out from the populated region to the adjacent regions. As a result, the spatial pattern gradually flattens and connects to $\bar{\bm{x}}$ at ϕsuperscriptitalic-$\phi^{**}$. We can interpret the region at the mode of population distribution (region $i$ such that $x_{i}>x_{i-1}$ and $x_{i}>x_{i+1}$ where mod $N$ for indices) as the location of an agglomeration. Then, this model endogenously produces at most one agglomeration.
@@
Type G model. Figure C.3 reports stable equilibrium patterns in the course of increasing $\phi$ for the Krugman-JPE1991 model. In Fig. 1C, the black solid (dashed) curves depict the stable (unstable) equilibrium values of $x_{i}$ at each $\phi$. Figure 1D is the schematic illustration of the stable spatial pattern on the path. The letters in Fig. 1D correspond to those in Fig. 1C. The global dispersion force in the Krugman model stems from competition between firms over consumers’ demand. If $\phi$ is low (if transport costs are high), firms have few incentives to agglomerate, and the uniform distribution is stable. If we increase $\phi$, competition with firms in other regions becomes fiercer, as the markets of other regions become closer. At some point firms are better off forming small agglomerations so that each agglomeration has its dominant market area but is relatively remote from other agglomerations of firms. At the so-called “break point” $\phi^{*}$, a bifurcation from $\bar{\bm{x}}$ occurs and the spatial pattern is pushed towards the formation of $\frac{8}{2}=4$ distinct agglomerations [Proposition 1 (b)]. A further increase in $\phi$ causes the second and third bifurcations at ϕsuperscriptitalic-$\phi^{**}$ and ϕsuperscriptitalic-$\phi^{***}$, respectively. These bifurcations sequentially double the spacing between agglomerations, each time halving their number, $4\to 2\to 1$, in a close analogy to the first bifurcation at $\phi^{*}$. We can formally analyze the successive bifurcations if we assume a specific model (ikeda2012spatial; Akamatsu-Takayama-Ikeda-JEDC2012; Osawa-et-al-JRS2017, see, e.g.,). At the higher extreme of $\phi$, agents concentrate in a single region. This behavior can be understood as a gradual extension of the market area of each agglomeration.
@@
Type LG model. With both local and global dispersion forces, Type LG models exhibit an interplay between the number of agglomerations, spacing between them (as in Type G models), and the spatial extent of each agglomeration (as in Type L models). Figure 2A shows the evolution of the number of agglomerations in the course of increasing $\phi$ under the Pfluger-Suedekum-JUE2008’s model (Section F.2.3). The number of agglomerations in a spatial distribution is defined by that of the local maxima therein. Figure 2A exhibits the mixed characteristics of Figs. C.3 and C.3, as expected. When $\phi<\phi^{*}$ or $\phi>\phi^{**}$, $\bar{\bm{x}}$ is stable. We interpret the number of agglomerations in $\bar{\bm{x}}$ as either $8$ (for a low $\phi$) or $1$ (for a high $\phi$) to acknowledge that $\bar{\bm{x}}$ at the low and high levels of $\phi$ are distinct. When $\phi$ gradually increases from $\phi\approx 0$, the number of agglomerations reduces from $8\to 4\to 2\to 1$ as in the Type G models (Fig. C.3), whereas it is always $1$ in the latter stage as per the Type L models (Fig. C.3). The initial stage is governed by a decline in the global dispersion force, while the later stage is marked by a relative rise of the local dispersion force.
@@ B Spatial pattern
Figure 2B illustrates the spatial patterns associated with Fig. 2A. Uniform pattern $\bar{\bm{x}}$ is initially stable (Pattern A) and the first bifurcation at $\phi^{*}$ leads to a quad-modal agglomeration (B, C), whereas the second bifurcation to the formation of a bimodal agglomeration (D, E). These transitions are in line with Fig. C.3 and are governed by the gradual decline in the global dispersion force. A further decline in the global dispersion force increases the relative importance of the local dispersion force. As a result, the bimodal agglomeration flattens out gradually (F, G). When $\phi$ increases further, it reduces to a unimodal agglomeration (J, K). The unimodal agglomeration flattens out as $\phi$ increases (L, M) until it converges to the complete dispersion (N) at ϕsuperscriptitalic-$\phi^{**}$.
Appendix D Geographical advantages
@@
The implications of Propositions 1, 2 and 3 qualitatively generalize to different geographical setups such as one-dimensional line segment, two-dimensional spaces with and without boundaries. In particular, the spatial distribution of agents become polycentric in Type G models, whereas it becomes monocentric in Type L models.
@@
The simplest way to introduce geographical asymmetry into our one-dimensional setup is to consider a bounded line segment, which is a standard stylized setting in urban economic theory. Ikeda-et-al-IJET2017 considered a Type G model (Forslid-Ottaviano-JoEG2003) in a line segment. They showed that multiple agglomerations emerge as in the circular economy and demonstrated that the evolution of spatial structure in a line segment approximately follows the “period doubling” behavior (Akamatsu-Takayama-Ikeda-JEDC2012; Osawa-et-al-JRS2017). For Types L and LG, Fig. D.1 reports examples of endogenous agglomeration patterns in the models by Helpman-Book1998 and Pfluger-Suedekum-JUE2008. For both models, qualitative properties of the spatial patterns are consistent with those discussed in Appendix C.
@@
The two-dimensional counterpart of the symmetric circle is bounded lattices with periodic boundary conditions, for which a basic theory of spatial agglomeration is provided in Ikeda-Murota-Book2014. For Type G models, they typically produce multiple disjointed agglomerations and period-doubling behavior as discussed in Appendix C (see, e.g., Ikeda-etal-IJBC2012; Ikeda-etal-JEBO2014; Ikeda-et-al-JRS2017; Ikeda-etal-JEDC2018). As concrete examples, Figure 10 in the main text shows endogenous equilibrium spatial patterns over a bounded square economy with $9\times 9=81$ regions in the course of increasing $\phi$ for the Krugman and Allen–Arkolakis models. The parameters are the same as Figs. C.3 and C.3. Their agglomeration processes are qualitatively consistent with Propositions 1, 2 and 3 and examples in Appendix C, suggesting the robustness of qualitative implications of our theoretical developments.
@@
The implications of Proposition 1 seem to extend to different assumptions on transport technology that are not formally covered by Assumption C. For example, linear transport costs are often assumed in the literature (e.g., Mossay-Picard-JET2011; Picard-Tabuchi-JET2013; Blanchet-et-al-IER2016). Mossay-Picard-JET2011 considered a variant of the Beckmann model (Type L) and showed that the only possible equilibrium is a unimodal distribution in a continuous line segment. Blanchet-et-al-IER2016 considered a general Type L model over a continuous one- or two-dimensional space; they showed that the equilibrium spatial pattern for the Beckmann model is unique and given by a regular concave paraboloid, i.e., a unimodal pattern. Picard-Tabuchi-JET2013 also considered a Type L general equilibrium model in a two-dimensional space and showed that spatial distribution becomes unimodal. The numerical results of anas1996general and anas1998urban in line segments bear a close resemblance to, respectively, agglomeration behaviors of Type L and G models, although they assume endogenous transport costs between locations.
Note: There are no asymmetries in regional characteristics except for geographical accessibility. The transport cost between every consecutive pair of regions is uniform. The level of transport cost monotonically decreases from top to bottom. Panels (A) and (B) consider a line segment with 65 locations. Panel (B) adapted from Ikeda-et-al-IJET2017 considers 17 locations; see the original paper for an extensive discussion on the behavior of a Type G model in a line segment.

### v8

@@ A.3  Proof of Proposition˜1
Proposition˜1 (a) implies that Type G models yield $\frac{N}{2}$ agglomerations, with further bifurcations of the form $\frac{N}{2}\to\frac{N}{4}\to\frac{N}{8}\to\cdots\to 2\to 1$ expected if $N$ is a power of two (Akamatsu-Takayama-Ikeda-JEDC2012; ikeda2012spatial; Osawa-et-al-JRS2017). For Type L, Takayama-Ikeda-Thisse-RSUE2020 confirmed the emergence of single-peaked or monocentric patterns in the (Murata-Thisse-JUE2005) model. AMT-2016 formally compares Forslid-Ottaviano-JoEG2003 (Type G) and Helpman-Book1998 (Type L). All available formal results in the literature corroborates with Proposition˜1 and the numerical examples in this study. ∎
A.4  Proof of Lemma˜4
@@
Under Assumption˜C, we can choose the same set of eigenvectors for $\mathbf{V}$ and $\mathbf{F}$ because they are both symmetric circulant matrices. Let $\{\bm{z}_{k}\}$ be the set of eigenvectors and let $\omega_{k}$ and $\eta_{k}$ be the eigenvalues of $\mathbf{V}$ and $\mathbf{F}$ associated with $\bm{z}_{k}$, respectively. Then, for each eigenvector $\bm{z}_{k}$ except for $\bm{z}_{0}=\bm{1}$, Eq.˜A.5 yields
@@
As $\mathbf{F}$ and $\mathbf{V}$ are both symmetric, $\eta_{k}$ and $\omega_{k}$ are both real. Thus, Eq.˜A.6 implies $\operatorname{sgn}[\eta_{k}]=\operatorname{sgn}[\omega_{k}]$. Therefore, $\bar{\bm{x}}$ is stable under all admissible dynamics if and only if $\omega_{k}<0$ for all $k$, excluding $k=0$ that corresponds to $\bm{z}_{0}=\bm{1}$. Likewise, $\bar{\bm{x}}$ is unstable if and only if $\omega_{k}>0$ for some $k$, again excluding $k=0$.
Suppose exactly one $\omega_{k}$ changes sign from negative to positive at $\phi_{k}^{*}$. Then, from Eq.˜A.6, the corresponding eigenvalue of the Jacobian matrix of any admissible dynamic at $\bar{\bm{x}}$ must also cross zero at $\phi_{k}^{*}$. Bifurcation theory shows that the system departs from $\bar{\bm{x}}$ along the direction of the associated eigenvector $\bm{z}_{k}$, as it is tangent to the “unstable manifold” at the bifurcation point (see, e.g., Hirsch-et-al-Book2012; Kuznetsov-Book2013).
@@ B.1  Development of high-speed transport networks in Japan
Figure˜B.1 reports the development of high-speed railway and highway networks, respectively, in Japan between 1970 and 2020. Both networks were initially spurred by infrastructure investments surrounding the 1964 Tokyo Olympics. Over this period, total highway length increased from 1,119 km to 9,050 km, while high-speed rail expanded from 515 km to 3,106 km, which are more than eightfold and sixfold increases, respectively. The steady, long-run expansion of these networks makes Japan a natural setting for studying comparative statics with respect to transport costs.
@@ B.2  Japanese cities and their growths
We identify cities in Japan using the Grid Square Statistics from the Population Census for 1970–2020. A city is defined as an urban agglomeration (UA), consisting of contiguous 1 km $\times$ 1 km grid cells with a population density of at least 1,000 persons/km2 and a total population of at least 10,000. Our results are robust to alternative threshold values. Figure˜B.2 displays the 431 UAs identified in 2020. These UAs occupy 6% of Japan’s land area while containing about 80% of the national population. Populated cells with fewer than 1,000 residents are shown in grey, with darker shading indicating higher population counts. We restrict the analysis to grid cells that are reachable by road from the four major islands—Hokkaido, Honshu, Shikoku, and Kyushu. UAs are identified separately for each census year from 1970 to 2020 (at five-year intervals), and consistent unique IDs are assigned across years to track individual agglomerations over time. For details on the construction of UAs, see Mori-Murakami-DP2025.
@@
From 1970 to 2020, the total population in these cities grew by $55\%$, while the national population increased by only 21%. The population of the largest city, Tokyo, has grown by 67%, an increase about the same size of the second largest city, Osaka. Figures˜B.3 and B.6 provide closer looks. Population growth is typically associated with concurrent areal growth (Fig.˜B.3A), but population density generally decreased during the 50 years period (Fig.˜B.3B), in particular for small to medium-sized cities, indicating local spreading of these cities. It is noted that the largest cities also experienced local spreading as evident in their spatial distributions shown in Fig.˜1C as well as Fig.˜B.4.
@@ B The population distribution within Nagoya in 1970 and 2020
Figure B.4: Local dispersion of Osaka (rank == 2) and Nagoya (rank == 3).
@@
Figure˜B.6 plots cities’ areal growth rates against their population growth rates, both measured on a logarithmic scale. By definition, $\log(\text{area})=\log(\text{population})-\log(\text{density})$, so the diagonal line in each panel represents constant population density (i.e., a density ratio of one). Cities above this line experienced a decline in density, whereas those below experienced an increase. The vertical and horizontal reference lines indicate Japan’s aggregate population growth ratio ($1.21$), and their intersection corresponds to a hypothetical city whose population and area both grew at the national average rate. For most cities, areal growth exceeded population growth. Relative to these reference lines, the northeast quadrant indicates simultaneous expansion in population and area (overall growth), while the southwest quadrant corresponds to joint decline, or urban shrinkage. The northwest quadrant reflects relative urban sprawl (area growth accompanied by relative population decline) whereas the southeast quadrant indicates relative densification (population growth accompanied by relatively slow area expansion). The southeast quadrant contains very few observations, suggesting that relative densification has been limited.
@@
Employment distribution. The tendency of local dispersion is observed in alternative indicator of agglomeration other than population. Figure˜B.6 shows the change in the mean values of the maximum and average employment density within a city across all cities in Japan from 1975 to 2014. Their long-run trend indicates that the geographical distribution of employment in a city has flattened over the past half century.
@@ B.3  Cities in other countries
Figure˜2 in Section˜1 uses Japanese Census data. To allow a parallel comparison between different countries, we employ the the LandScan™ Global Population Database, developed by the Department of Energy’s Oak Ridge National Laboratory (ORNL), as the basic grid population data to see the evolution of cities. Cities are defined in the same manner as for Japanese cities. To asses compatibility between the LandScan data and the census we compare Census and the LandScan data for Japan. Figures˜7A and 7B are, respectively, based on the Census and the the LandScan data. The LandScan data is only available after 2000, and are only roughly in agreement with the precise data based on Census. It is also noted that, while the LandScan data is available every year, we employ 5-year steps to avoid noises due to its data generation procedure that incorporates various interpolations. Nonetheless, Figure˜B.7 confirms the broad tendency of nationwide concentration and local flattening of the cities. To be consistent with our theory that assume a fixed total population, we normalize the total population in each country to unity. In the cases of France, Germany and Japan, the local flattening is apparent even without the normalization.
A Japan (1970–2020), Census
[…]
Figure B.7: Global concentration and local dispersion
@@ Appendix C Eight regions
This section considers agglomeration processes in the $N=8$ circular economy. For a selected model from each model category, we follow stationary equilibria branching from $\bar{\bm{x}}$ and then numerically check the local stability of those stationary equilibrium solutions under the replicator dynamic (Taylor-Jorker-MB1978).
@@
Type L model. Figure˜C.3 considers a Type L model by Allen-Arkolakis-QJE2014 (Section˜F.2.4; we set $\alpha=0.5$, $\beta=-0.3$, and $\sigma=6.0$). The model incorporates a local dispersion force but no global dispersion force. The uniform equilibrium $\bar{\bm{x}}$ is stable when transport costs are low (when $\phi$ is close to $1$). If we start from $\bar{\bm{x}}$ and consider the process of a monotonic decrease in $\phi$ from $\phi\approx 1$, then a unimodal pattern emerges due to the bifurcation at $\phi^{**}$ [Proposition˜1 (a)]. This is the bifurcation in the model. When $\phi$ decreases further, the spatial pattern smoothly converges to a full concentration in a single region in the lower extreme ($\phi\approx 0$). The local dispersion force is less important than the benefits of agglomeration when interregional transportation is prohibitively costly. Mobile agents prefer concentrating on a smaller number of regions because of the agglomeration forces. As $\phi$ increases, agglomeration force due to costly transportation diminishes, and the relative rise in the local dispersion force induces a crowding-out from the populated region to the adjacent regions. As a result, the spatial pattern gradually flattens and connects to $\bar{\bm{x}}$ at $\phi^{**}$. We can interpret the region at the mode of population distribution (region $i$ such that $x_{i}>x_{i-1}$ and $x_{i}>x_{i+1}$ where mod $N$ for indices) as the location of an agglomeration. Then, this model endogenously produces at most one agglomeration.
@@
Type G model. Figure˜C.3 reports stable equilibrium patterns in the course of increasing $\phi$ for the Krugman-JPE1991 model (Section˜F.2.1; we set $\mu=0.5$, $\sigma=10$, and $L=8$.). In Fig.˜1C, the black solid (dashed) curves depict the stable (unstable) equilibrium values of $x_{i}$ at each $\phi$. Figure˜1D is the schematic illustration of the stable spatial pattern on the path. The letters in Fig.˜1D correspond to those in Fig.˜1C. The global dispersion force in the Krugman model stems from competition between firms over consumers’ demand. If $\phi$ is low (if transport costs are high), firms have few incentives to agglomerate, and the uniform distribution is stable. If we increase $\phi$, competition with firms in other regions becomes fiercer, as the markets of other regions become closer. At some point firms are better off forming small agglomerations so that each agglomeration has its dominant market area but is relatively remote from other agglomerations of firms. At the so-called “break point” $\phi^{*}$, a bifurcation from $\bar{\bm{x}}$ occurs and the spatial pattern is pushed towards the formation of $\frac{8}{2}=4$ distinct agglomerations [Proposition˜1 (b)]. A further increase in $\phi$ causes the second and third bifurcations at $\phi^{**}$ and $\phi^{***}$, respectively. These bifurcations sequentially double the spacing between agglomerations, each time halving their number, $4\to 2\to 1$, in a close analogy to the first bifurcation at $\phi^{*}$. We can formally analyze the successive bifurcations if we assume a specific model (ikeda2012spatial; Akamatsu-Takayama-Ikeda-JEDC2012; Osawa-et-al-JRS2017, see, e.g.,). At the higher extreme of $\phi$, agents concentrate in a single region. This behavior can be understood as a gradual extension of the market area of each agglomeration.
@@
Type LG model. With both local and global dispersion forces, Type LG models exhibit an interplay between the number of agglomerations, spacing between them (as in Type G models), and the spatial extent of each agglomeration (as in Type L models). Figure˜2A shows the evolution of the number of agglomerations in the course of increasing $\phi$ under the Pfluger-Suedekum-JUE2008’s model (Section˜F.2.3; we set $\mu=0.4$, $\sigma=2.5$, $L=4$, $\gamma=0.5$, and $a_{i}=1$). The number of agglomerations in a spatial distribution is defined by that of the local maxima therein. Figure˜2A exhibits the mixed characteristics of Figs.˜C.3 and C.3, as expected. When $\phi<\phi^{*}$ or $\phi>\phi^{**}$, $\bar{\bm{x}}$ is stable. We interpret the number of agglomerations in $\bar{\bm{x}}$ as either $8$ (for a low $\phi$) or $1$ (for a high $\phi$) to acknowledge that $\bar{\bm{x}}$ at the low and high levels of $\phi$ are distinct. When $\phi$ gradually increases from $\phi\approx 0$, the number of agglomerations reduces from $8\to 4\to 2\to 1$ as in the Type G models (Fig.˜C.3), whereas it is always $1$ in the latter stage as per the Type L models (Fig.˜C.3). The initial stage is governed by a decline in the global dispersion force, while the later stage is marked by a relative rise of the local dispersion force.
@@ B Spatial pattern
Figure˜2B illustrates the spatial patterns associated with Fig.˜2A. Uniform pattern $\bar{\bm{x}}$ is initially stable (Pattern A) and the first bifurcation at $\phi^{*}$ leads to a quad-modal agglomeration (B, C), whereas the second bifurcation to the formation of a bimodal agglomeration (D, E). These transitions are in line with Fig.˜C.3 and are governed by the gradual decline in the global dispersion force. A further decline in the global dispersion force increases the relative importance of the local dispersion force. As a result, the bimodal agglomeration flattens out gradually (F, G). When $\phi$ increases further, it reduces to a unimodal agglomeration (J, K). The unimodal agglomeration flattens out as $\phi$ increases (L, M) until it converges to the complete dispersion (N) at $\phi^{**}$.
Appendix D Geographic advantages
@@
The implications of Propositions˜1, 2 and 3 qualitatively generalize to different settings such as one-dimensional line segment, two-dimensional spaces. The spatial distribution of agents becomes polycentric in Type G models, whereas it becomes monocentric in Type L models.
@@
The simplest way to introduce geographic asymmetry into our one-dimensional setting is to consider a bounded line segment, which is a standard stylized setting in urban economic theory. Ikeda-et-al-IJET2017 considered a Type G model (Forslid-Ottaviano-JoEG2003) in a line segment. They showed that multiple agglomerations emerge as in the circular economy and demonstrated that the evolution of spatial structure in a line segment approximately follows the “period doubling” behavior (Akamatsu-Takayama-Ikeda-JEDC2012; Osawa-et-al-JRS2017). For Types L and LG, Fig.˜D.1 reports examples of endogenous agglomeration patterns in the models by Helpman-Book1998 and Pfluger-Suedekum-JUE2008. For both models, qualitative properties of the spatial patterns are consistent with those discussed in Appendix˜C.
@@
The two-dimensional counterpart of the symmetric circle is bounded lattices with periodic boundary conditions, for which a basic theory of spatial agglomeration is provided in Ikeda-Murota-Book2014. For Type G models, they typically produce multiple disjointed agglomerations and period-doubling behavior as discussed in Appendix˜C (see, e.g., Ikeda-etal-IJBC2012; Ikeda-etal-JEBO2014; Ikeda-et-al-JRS2017; Ikeda-etal-JEDC2018). As concrete examples, Figure˜11 in the main text shows endogenous equilibrium spatial patterns over a bounded square economy with $9\times 9=81$ regions in the course of increasing $\phi$ for the Krugman and Allen–Arkolakis models. The parameters are the same as Figs.˜C.3 and C.3. Their agglomeration processes are qualitatively consistent with Propositions˜1, 2 and 3 and examples in Appendix˜C, suggesting the robustness of qualitative implications of our theoretical developments.
@@
The implications of Proposition˜1 seem to extend to different assumptions on transport technology that are not formally covered by Assumption˜C. For example, linear transport costs are often assumed in the literature (e.g., Mossay-Picard-JET2011; Picard-Tabuchi-JET2013; Blanchet-et-al-IER2016). Mossay-Picard-JET2011 considered a variant of the Beckmann model (Type L) and showed that the only possible equilibrium is a unimodal distribution in a continuous line segment. Blanchet-et-al-IER2016 considered a general Type L model over a continuous one- or two-dimensional space; they showed that the equilibrium spatial pattern for the Beckmann model is unique and given by a regular concave paraboloid, i.e., a unimodal pattern. Picard-Tabuchi-JET2013 also considered a Type L general equilibrium model in a two-dimensional space and showed that spatial distribution becomes unimodal. The numerical results of anas1996general and anas1998urban in line segments bear a close resemblance to, respectively, agglomeration behaviors of Type L and G models, although they assume endogenous transport costs between locations.